a box contains 8 red,3 white and 9 blue balls if three balls are drawn without replacement determine the probability
a)all three are red
b)two are red and one is white
c)all three are white
d)at least one is white
e)one for each color is drawn
1
Expert's answer
2012-09-13T11:00:51-0400
There are 8+3+9 = 20 balls in a box.
p{r} = 8/20 = 0.4;
p{w} = 3/20 = 0.15;
p{b} = 9/20 = 0.45;

a. P = 8/20 * 7/19 * 6/18 = 336/6840 = 14/285 ( = C(8,3)/C(20,3) );
b. P = (C(8,2) + C(3,1)) / C(20,3) = 31/1140;
c. P = 3/20 * 2/19 * 1/18 = 6/6840 = 1/1140 ( = C(3,3)/C(20,3) );
d. P = (C(3,1) + C(19,2))/C(20,3) = 174/1140 = 87/570;
e. P = p{r}p{w}p{b} = 0.4*0.15*0.45 = 0.027.